# Hall Viscosity in the Quark-Gluon Plasma

### 24 Mar 2026 v1

@@
2 Department of Physics Gleb Watagin, University of Campinas, Campinas, Brazil
@@ 3.2 Observables and their quantitative estimates
Additional dynamical fluctuations can be constrained by $v_{2}$ fluctuation and correlations, in other words event-by-event [75, 76, 77, 78]
$\left\langle\left(\delta v_{n}\delta v_{m}\right)\right\rangle\equiv\langle\cos[n(\phi-\phi_{0n})]\cos[m(\phi-\phi_{0m})]\rangle-\langle\cos[n(\phi-\phi_{0n})]\rangle\langle\cos[m(\phi-\phi_{0m})]\rangle\,.$ (57)
Considering the odd components, correlations between $\phi_{n}$ and $\phi_{m}$ can also be isolated. Events can even be binned by $\cos[n(\phi-\phi_{0n})]$ leading to “event engineering” [79] (geometry selection) techniques. So far no evidence of fluctuations due to dynamics have been found, see [80] and references therein.
From the previous sections, it is straightforward to see that Hall viscosity appears in this latter category. In particular, the longitudinal component of vorticity is directly connected to the well-developed program of the study of fluctuations of anisotropic flow/event eccentricity. [^4] The transverse component of vorticity, which dominates at lower energy, could also lead to interesting effects in rapidity, but the background of these fluctuations is not as well understood. For both longitudinal and transverse Hall viscosity, a distinctive signature is the correlation between
$\theta_{Q}=\mathrm{Arg}[Q]\,,\qquad Q=p_{x}+ip_{y}$
and the global geometry (impact parameter distribution).
@@ 4 Conclusions and Outlook
Incorporating $\tilde{\eta}_{\perp}$ and $\tilde{\eta}_{\parallel}$ into relativistic magneto-hydrodynamic codes will allow a quantitative study of how Hall viscous terms modify freeze-out observables. [^5] For a self-consistent (magneto-)hydrodynamic description, also the other anisotropic transport coefficients should be included, most prominently the longitudinal and transverse shear viscosities, as well as the longitudinal and transverse electric conductivities [88, 89, 90, 91, 92, 93, 94, 64, 62, 63]. Generically, with increasing anisotropy (stronger magnetic field, stronger rotation) the longitudinal transport coefficient will deviate more from the transverse one. The predicted phase-angle shifts $(\Delta\theta,\Delta\phi)$ could be compared to event-plane correlations and $\Lambda$-hyperon global polarization data. The coupling of shear to spin polarization [95, 96, 97] suggests a direct relation between Hall viscosity and spin hydrodynamics that warrants investigation. Future global fits could include $\tilde{\eta}_{\perp}$ and $\tilde{\eta}_{\parallel}$ as additional transport parameters in Bayesian analyses [44, 98, 99, 100]. Sensitivity studies using observables such as elliptic flow $v_{2}$, directed flow $v_{1}$, polarization correlations, or event-plane tilts would constrain the magnitude and sign of Hall viscosities in the QGP.
@@ Acknowledgements
We thank Dmitri Kharzeev for valuable comments and Aleksi Kurkela for discussions in the early stage of this project. R.M. acknowledges the support of the German Research Foundation (DFG) through the Collaborative Research Center ToCoTronics, Project-ID 258499086 — SFB 1170, as well as Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ctd.qmat (EXC 2147, Project-ID 390858490). He furthermore acknowledges hospitality from the Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS) during the final stages of this work, and associated travel support under STCSM Grant 25HB2701900. G.T. thanks Bolsa de produtividade CNPQ 305731/2023-8 and FAPESP 2023/06278-2, as well as participation in the tematico 2023/13749-1 for support. M.K. was supported, in part, by the U.S. Department of Energy grant DE-SC0012447. M.K. thanks the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work. S.M. thanks the Theoretical Physics III Group at the University of Würzburg for their hospitality and support during the early stages of this work.

### 30 Mar 2026 v2

@@
2 Instituto de Fisica Gleb Wataghin, University of Campinas, Campinas, Brazil
@@ 3.2 Observables and their quantitative estimates
From the previous sections, it is straightforward to see that Hall viscosity would result in an event plane and tilt plane rotations with respect to their ”geometric” expecations. In particular, the longitudinal component of vorticity is directly connected to the well-developed program of the study of fluctuations of anisotropic flow/event eccentricity. [^4] The transverse component of vorticity, which dominates at lower energy, could also lead to interesting effects in rapidity, but the background of these fluctuations is not as well understood. We shall briefly discuss experimental observables sensitive to this, as well as make a few qualitative estimates.
Dynamical fluctuations can be constrained by $v_{2}$ fluctuation and correlations, expressible event-by-event via symmetric cumulants[75, 76, 77, 78, 79] ($\Delta A=A-\left\langle A\right\rangle$)
$\left\langle\left(\delta v_{n}(y)\delta v_{m}(y^{\prime})\right)\right\rangle\equiv\langle\Delta\cos[n(\phi-\phi_{0n}(y))]\Delta\cos[m(\phi-\phi_{0m}(y^{\prime})]\rangle_{event}$ (57)
$\sim\left\langle\Delta\cos\left(n\left(\sum_{i}^{M}\phi_{i}-\sum_{j}^{M}\phi_{j}\right)\right)\Delta\cos\left(m\left(\sum_{i}^{K}\phi_{k}-\sum_{j}^{K}\phi_{l}\right)\right)\right\rangle_{event}$
With a separation between y and y’ [80]. Higher order cumulants $M,K\gg 2$ and greater rapidity separation eliminats ”non-flow” effects (jet fragmentation, resonance decays, momentum conservation and so on). Hall viscosity is unusual in that it is a ”flow” effect (and hence will persist for higher symmetric cumulants) but the rotation of $\phi_{0}$ will be local in rapidity. Provided the detector has enough rapidity coverage and enough statistics to check cumulant scaling, the effect of Hall viscosity might qualitatively appear.
Correlations between $\phi_{n}$ and $\phi_{m}$ can also be isolated (if $y\rightarrow y^{\prime}$ the cumulant is by construction independent of $\phi_{0}$). More direct observables of $\phi_{0}$ are also possible [81, 82] and, provided they can be done at wide rapidity separations (or an event plane detector is available) , they can directly infer a rotation of the azimuthal momentum event plane w.r.t. its geometric expectation.
Events can even be binned by $\cos[n(\phi-\phi_{0n})]$ leading to “event engineering” [83] (geometry selection) techniques. So far no evidence of fluctuations due to dynamics have been found, see [84] and references therein.
@@ 4 Conclusions and Outlook
Incorporating $\tilde{\eta}_{\perp}$ and $\tilde{\eta}_{\parallel}$ into relativistic magneto-hydrodynamic codes will allow a quantitative study of how Hall viscous terms modify freeze-out observables. [^5] For a self-consistent (magneto-)hydrodynamic description, also the other anisotropic transport coefficients should be included, most prominently the longitudinal and transverse shear viscosities, as well as the longitudinal and transverse electric conductivities [92, 93, 94, 95, 96, 97, 98, 64, 62, 63]. Generically, with increasing anisotropy (stronger magnetic field, stronger rotation) the longitudinal transport coefficient will deviate more from the transverse one. The predicted phase-angle shifts $(\Delta\theta,\Delta\phi)$ could be compared to event-plane correlations such as Eq. (57), more direct probes of event plane geometry [81, 82] and $\Lambda$-hyperon global polarization data. The coupling of shear to spin polarization [99, 100, 101] suggests a direct relation between Hall viscosity and spin hydrodynamics that warrants investigation. Future global fits could include $\tilde{\eta}_{\perp}$ and $\tilde{\eta}_{\parallel}$ as additional transport parameters in Bayesian analyses [44, 102, 103, 104]. Sensitivity studies using observables such as elliptic flow $v_{2}$, directed flow $v_{1}$, polarization correlations, or event-plane tilts would constrain the magnitude and sign of Hall viscosities in the QGP.
@@ Acknowledgements
We thank Dmitri Kharzeev for valuable comments and Aleksi Kurkela for discussions in the early stage of this project, as well as Ante Biladzic and Jiangyong Jia for helpful comments and suggestions on the manuscript. R.M. acknowledges the support of the German Research Foundation (DFG) through the Collaborative Research Center ToCoTronics, Project-ID 258499086 — SFB 1170, as well as Germany’s Excellence Strategy through the Würzburg-Dresden Cluster of Excellence on Complexity and Topology in Quantum Matter - ctd.qmat (EXC 2147, Project-ID 390858490). He furthermore acknowledges hospitality from the Shanghai Institute for Mathematics and Interdisciplinary Sciences (SIMIS) during the final stages of this work, and associated travel support under STCSM Grant 25HB2701900. G.T. thanks Bolsa de produtividade CNPQ 305731/2023-8 and FAPESP 2023/06278-2, as well as participation in the tematico 2023/13749-1 for support. M.K. was supported, in part, by the U.S. Department of Energy grant DE-SC0012447. M.K. thanks the Galileo Galilei Institute for Theoretical Physics for the hospitality and the INFN for partial support during the completion of this work. S.M. thanks the Theoretical Physics III Group at the University of Würzburg for their hospitality and support during the early stages of this work.
@@
  • [79] J. Adam et al., Correlated event-by-event fluctuations of flow harmonics in Pb-Pb collisions at $\sqrt{s_{{}_{\rm NN}}}=2.76$ TeV, Phys. Rev. Lett. 117, 182301 (2016), 10.1103/PhysRevLett.117.182301, 1604.07663.
  • [80] C. Mordasini, A. Bilandzic, D. Karakoç and S. F. Taghavi, Higher order Symmetric Cumulants, Phys. Rev. C 102(2), 024907 (2020), 10.1103/PhysRevC.102.024907, 1901.06968.
  • [81] S. Acharya et al., Symmetry plane correlations in Pb-Pb collisions at $\sqrt{s_{\rm NN}}=2.76$TeV, Eur. Phys. J. C 83(7), 576 (2023), 10.1140/epjc/s10052-023-11658-w, 2302.01234.
  • [82] G. Aad et al., Measurement of event-plane correlations in $\sqrt{s_{NN}}=2.76$ TeV lead-lead collisions with the ATLAS detector, Phys. Rev. C 90(2), 024905 (2014), 10.1103/PhysRevC.90.024905, 1403.0489.